OpenStudy (anonymous):

solve for x 4(3x-2)=-2(x+4)

OpenStudy (anonymous):

4(3x-2)/-2=-2(x+4)/-2 -6x+4=x+4 7x=0 x=0

OpenStudy (anonymous):

12x-8=-2x-8 14x=0 x=0